adding and subtracting rational expressions

(x+2)/4 + (3x+1)/3
1
Expert's answer
2012-12-07T03:09:33-0500
(x+2)/4 + (3x+1)/3={3(x+2)+4(3x+1)}/12={3x+6+12x+4}/12={15x+10}/12=15x/12+10/12=5x/4+5/6